3) A manufacturer would like to determine if the variance in a product dimension exceeds 10. a) State the Hypothesis to show the variance is greater than 10. b) Choose a level of a Use a= 0.05 for this problem. c) To test the hypothesis, the manufacturer takes a sample of 18 parts and measure that product dimension. The data appear in the Dimension worksheet of the HW3 data workbook on Moodle. Collect data and calculate necessary statistics to test the hypothesis. d) Sketch the sampling distribution. Include the critical value and test statistic. e) Draw a conclusion and report that in the problem context. f) What is the p-value for the hypothesis test?

Part Data
1 48.3
2 48.5
3 54.0
4 49.3
5 48.8
6 56.6
7 45.8
8 55.2
9 45.0
10 52.0
11 46.4
12 51.2
13 50.4
14 47.6
15 45.1
16 44.8
17 44.9
18 56.3
Answer #1

Solution:

Given: A manufacturer would like to determine if the variance in a product dimension exceeds 10.

Sample size = n = 18

Part a) State the Hypothesis to show the variance is greater than 10.

Ho: o210

Vs

H1 02> 10 :

Part b) Choose a level of a Use a= 0.05 for this problem

Find critical value for a significance level = 0.05

df = n - 1 = 18 - 1 = 17

x 00 x950 x975 x900 X990 df 050 995 3.841 5.991 7,815 9.488 11.070 0,001 1 0.000 0.000 0.004 0.016 2.706 0.020 2 0.010 0.051

Chi-square critical value = 27.587

27.587 Xcrit

Part c) Find sample variance and Chi-square test statistic for variance.

(n 1) x s 2

where

s^{2}= \frac{\sum x^{2}-(\sum x)^{2}/n}{n-1}

Thus we need to make following table:

Part Data x x 2
1 48.3 2332.89
2 48.5 2352.25
3 54.0 2916
4 49.3 2430.49
5 48.8 2381.44
6 56.6 3203.56
7 45.8 2097.64
8 55.2 3047.04
9 45.0 2025
10 52.0 2704
11 46.4 2152.96
12 51.2 2621.44
13 50.4 2540.16
14 47.6 2265.76
15 45.1 2034.01
16 44.8 2007.04
17 44.9 2016.01
18 56.3 3169.69
890.2 244297.38

Thus we get:

s^{2}= \frac{\sum x^{2}-(\sum x)^{2}/n}{n-1}

44297.38 (890.2)2/18 2 18 1.

44297.38 44025.33556 2 17

272.04444 2 17

s^{2}=16.00261

Thus we get:

(n 1) x s 2

(18 1) x 16.00261 2 10

17 x 16.00261 2 10

\chi ^{2}=27.204

Part  d) Sketch the sampling distribution. Include the critical value and test statistic.

Rejection Region 0.05 227.204 Xerit 2 27.587

Part e) Draw a conclusion and report that in the problem context.

Decision rule: Reject null hypothesis H0, if Chi-square test statistic value > Chi-square critical value = 27.587, otherwise we fail to reject H0.

Since Chi-square test statistic value =\chi ^{2}=27.204 < Chi-square critical value = 27.587,we fail to reject H0.

Thus there is not sufficient evidence to conclude that: the variance in a product dimension exceeds 10.

Part f) What is the p-value for the hypothesis test?

To get exact p-value, use Excel command:

=CHISQ.DIST.RT( x2 , df )

=CHISQ.DIST.RT(27.204 , 17 )

=0.0551

p-value = 0.0551

To get interval of p-value using table, look in Chi-square table for df = 17 row and find interval in which \chi ^{2}=27.204 fall, then find corresponding right tail area interval , which would be range of p-value.

x995 x373 x950 x300 xoso df .990 100 .050 .025 3841 5991 7815 9 488 11.070 12 592 14 067 15507 16919 18 307 19 675 21 026 22.

\chi ^{2}=27.204 fall between 24.769 and 27.587 , corresponding right tail area is between 0.05 to 0.10

Thus 0.05 < p-value < 0.10